# Attribute Inference Attacks for Federated Regression Tasks

### v1

@@ 6 Discussion and conclusions
Last but not least, to mitigate privacy leakage, beyond differential privacy mechanisms, there are empirical defenses such as Mixup (Zhang et al. 2018), which trains a model on composite data samples obtained through linear combinations of sample pairs, and TAPPFL (Arroyo Arevalo et al. 2024), which learns low-dimensional representations that minimize information about private attributes. However, the effectiveness of these defenses has not yet been demonstrated on regression tasks, and we leave this for future work.
7 Acknowledgements
This research was supported in part by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge, as well as by the France 2030 program, managed by the French National Research Agency under grant agreements No. ANR-21-PRRD-0005-01, ANR-23-PECL-0003, and ANR-22-PEFT-0002. It was also funded in part by the European Network of Excellence dAIEDGE under Grant Agreement Nr. 101120726, by SmartNet and LearnNet, and by the French government National Research Agency (ANR) through the UCA JEDI (ANR-15-IDEX-01), EUR DS4H (ANR-17-EURE-004), and the 3IA Côte d’Azur Investments in the Future project with the reference number ANR-19-P3IA-0002.
@@ A.1 Full statement and proof of Theorem 1
Consider training a least squares regression through FedAvg (Alg. 2) with batch size $B$ and local epochs $E$. Assume that
@@
    the client’s design matrix $\mathbf{x}_{c}\in\mathbb{R}^{m\times d}$ has rank $d$ equal to the number of features plus one; [^5] Remember that a client’s design matrix has a number of rows equal to the input features of the samples in the client’s local training dataset.
@@
    the components of the stochastic (mini-batch) gradient are distributed as sub-Gaussian random variables with variance proxy $\sigma^{2}$, i.e., $\mathbb{E}[\exp(\mathbf{e}(\theta)[i]/\sigma^{2})]\leq\exp(1),\forall\theta,\forall i\in\{1,2,...,d\},$ where $\mathbf{e}(\theta)=\mathbf{g}(\theta)-\nabla{\mathcal{L}_{c}}(\theta)$; [^6] Note that, by Jensen’s inequality, this condition implies a bounded variance for each component of the stochastic gradients.
@@
    there exists $\underline{\lambda}>0$ such that $\forall n_{c}\in\mathbb{N}$, we can always select $n_{c}$ observation rounds so that $\lambda_{\min}\!\left(\frac{\Theta_{\text{out}}^{T}\Theta_{\text{out}}}{n_{c}}\right)\geq\underline{\lambda}$, where $\lambda_{\min}(\mathbf{A})$ denotes the smallest eigenvalue of the matrix $\mathbf{A}$, and $\Theta_{\text{out}}$ is defined in Alg. 3.
The error of the reconstructed model $\hat{\theta}^{*}_{c}$ of Algorithm 3 is upper bounded w.p. $\geq 1-\delta$ when $\eta\leq\frac{m}{2\lambda_{\max}(\mathbf{x}_{c}^{T}\mathbf{x}_{c})}$ and
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}=\mathcal{O}\left(\eta\sigma d\sqrt{dE\left\lceil\frac{m}{B}\right\rceil\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\cdot\underline{\lambda}}}\right).$ (7)
@@
Let $\mathbf{H}=\mathbf{x}_{c}^{T}\mathbf{x}_{c}$ which is positive definite. Let $\mathbf{y}_{c}\in\mathbb{R}^{m}$ be the labels in the local dataset ${\mathcal{D}}_{c}$ with size $m=|{\mathcal{D}}_{c}|$.
${\mathcal{L}_{c}}(\theta)=\frac{\left\lVert\mathbf{x}_{c}\theta-\mathbf{y}_{c}\right\rVert^{2}}{m}$ (8)
@@
$\mathbf{g}(\theta)=\nabla{\mathcal{L}_{c}}(\theta)+\mathbf{e}(\theta)=\frac{2}{m}\left(\mathbf{H}\theta-\mathbf{H}\theta_{c}^{*}\right)+\mathbf{e}(\theta).$ (9)
@@
$\theta_{c}^{t}(K)$ $=\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{K}\theta_{c}^{t}(0)+\left[\mathbf{I}-\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{K}\right]\theta_{c}^{*}$
$-\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{k-1}\eta\mathbf{e}_{K-k+1}.$
where $K=E\lceil\frac{m}{B}\rceil$ as in each epoch there are $\lceil\frac{m}{B}\rceil$ local steps. Let $\mathbf{W}=\left[\mathbf{I}-\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{K}\right]$. $\mathbf{W}$ is proven to be invertible in Lemma 2, then we have:
@@
$-\mathbf{W}^{-1}\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{k-1}\eta\mathbf{e}_{K-k+1}$ (10)
Let $\mathbf{\gamma}^{t}=-\eta\mathbf{W}^{-1}\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{k-1}\mathbf{e}_{K-k+1}$. Let $\mathbf{x}[i]$ be the $i^{th}$ element of vector $\mathbf{x}$ and $\mathbf{X}[i,:]$ and $\mathbf{X}[:,i]$ be the $i^{th}$ row and column of the matrix $\mathbf{X}$, respectively. Since every element in $\mathbf{e}$ is sub-Gaussian and the weighted sum of independent sub-Gaussian variables is still sub-Gaussian, $\mathbf{\gamma}[i]$ is sub-Gaussian with $\mathbb{E}[\mathbf{\gamma}^{t}[i]]=0$ and
$\mathbb{V}\mathrm{ar}[\mathbf{\gamma}^{t}[i]]$ $\leq d\eta^{2}\sigma^{2}\lVert\mathbf{W}^{-1}\rVert_{2}^{2}\sum_{k=1}^{K}\left\lVert\left(\mathbf{I}-\frac{2\eta}{m}\mathbf{H}\right)^{k-1}\right\rVert^{2}_{2}$
$=\frac{d\eta^{2}\sigma^{2}}{\lambda^{2}_{\text{min}}(\mathbf{W})}\sum_{k=1}^{K}\left\lVert(\mathbf{I}-\frac{2\eta}{m}\mathbf{H})^{k-1}\right\rVert^{2}_{2}$
$=\frac{d\eta^{2}\sigma^{2}}{\lambda^{2}_{\text{min}}(\mathbf{W})}\sum_{k=1}^{K}\rho_{\max}^{2}((\mathbf{I}-\frac{2\eta}{m}\mathbf{H})^{k-1}),$
@@
Let $\Lambda$ be a diagonal matrix whose entries are the positive eigenvalues of $\mathbf{H}$ and $\lambda_{\max}(\mathbf{H})$ be the largest eigenvalue of $\mathbf{H}$, i.e., $\lambda_{\max}(\mathbf{H})=\lVert\Lambda\rVert_{1}$. When $\eta\leq\frac{m}{2\lambda_{\max}(\mathbf{H})}$, the diagonal values in $\frac{2\eta}{m}\Lambda$ are positive and smaller than or equal to $1$. Moreover, the matrix $(\mathbf{I}-\frac{2\eta}{m}\mathbf{H})^{k-1}$ is positive semi-definite and we have $\rho_{\max}[(\mathbf{I}-\frac{2\eta}{m}\mathbf{H})^{k-1}]\leq 1$. Thus, we have
@@
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}^{2}\leq\frac{d^{2}\eta^{2}\sigma^{2}K}{\lambda_{\text{min}}^{2}(\mathbf{W})}\cdot\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\underline{\lambda}},\;\;\text{w.p.}\geq 1-\delta.$ (18)
@@
$\mathbf{W}$ $=I-\left(I-\frac{2\eta}{m}\mathbf{H}\right)^{E\left\lceil\frac{m}{b}\right\rceil}$
$=\mathbf{U}\left(I-\left(I-\frac{2\eta}{m}\Lambda\right)^{E\left\lceil\frac{m}{b}\right\rceil}\right)\mathbf{U}^{T}.$
@@
$\lambda_{\min}(\mathbf{W})$ $=1-\left(1-\frac{2\eta}{m}\lambda_{\min}(\mathbf{H})\right)^{E\left\lceil\frac{m}{b}\right\rceil}$
$\lambda_{\max}(\mathbf{W})$ $=1-\left(1-\frac{2\eta}{m}\lambda_{\max}(\mathbf{H})\right)^{E\left\lceil\frac{m}{b}\right\rceil}$
@@ A.2 Proof of Proposition 2
$\mathbf{g}(\theta)=\nabla{\mathcal{L}_{c}}(\theta)=\frac{2}{m}\left(\mathbf{H}\theta-\mathbf{H}\theta_{c}^{*}\right)$
@@ A.3 Proof of Proposition 3
${\mathcal{L}_{\mathbf{c}}}(\theta)$ $=$ $\frac{1}{m}\left(\theta^{T}\mathbf{H}_{\mathbf{c}}\theta-2(\mathbf{x}_{c}^{T}\mathbf{y}_{c})^{T}\theta-\mathbf{y}_{c}^{T}\mathbf{y}_{c}\right)$ (20)
$=$ $\frac{1}{m}\left(\sum_{i=1}^{d}\theta_{[i]}^{2}-\sum_{i=1}^{d-1}\theta_{[i]}\theta_{[i+1]}-\theta_{[1]}-C\right)$
@@
Let $n=|{\mathcal{C}}|$ be the number of clients and assume that every client has $m$ local data samples. Then the global empirical risk and its gradients have the following expressions:
${\mathcal{L}}(\theta)$ $=\frac{1}{nm}\left(\sum_{i=1}^{d}n\times\theta_{[i]}^{2}-\sum_{i=1}^{d-1}\theta_{[i]}\theta_{[i+1]}-\theta_{[1]}-C^{\prime}\right),$ (21)
$\nabla{\mathcal{L}}(\theta)_{[1]}$ $=\frac{1}{nm}(2n\times\theta_{[1]}-\theta_{[2]}-1),$ (22)
$\nabla{\mathcal{L}}(\theta)_{[i]}$ $=\frac{1}{nm}\left(2n\times\theta_{[i]}-\theta_{[i-1]}-\theta_{[i+1]}\right),\forall i\in\{2,...,d-1\},$ (23)
$\nabla{\mathcal{L}}(\theta)_{[d]}$ $=\frac{1}{nm}\left(2n\times\theta_{[d]}-\theta_{[d-1]}\right),$ (24)
@@ Appendix B The details of experimental setups
B.1 Experiment in Figure 1
We evaluate different AIAs when four clients train a neural network (a single hidden layer of 128 neurons using ReLU as activation function) through FedAvg with 1 local epoch and batch size 32. Each client stores $|{\mathcal{D}}_{c}|$ data points randomly sampled from ACS Income dataset (Ding et al. 2024). This dataset contains census information from 50 U.S. states and Puerto Rico, spanning from 2014 to 2018. It includes 15 features related to demographic information such as age, occupation, and education level. The adversary infers the gender attribute of every data sample held by the client given access to the released (public) information.
@@ B.1 Experiment in Figure 1
B.2 Toy dataset setting
We test our proposed passive LMRA (Alg. 3) on a toy federated least squares regression with two clients. Each client $c$ has 1024 samples with 10 features, where nine numerical features are sampled from a uniform distribution over $[0,1)$ and one binary feature is sampled from a Bernoulli distribution with $p=1/2$. The prediction $\mathbf{y}$ is generated from the regression model $\mathbf{y}=\mathbf{x}_{c}\theta^{*}+\mathbf{\epsilon}$ where $\mathbf{\epsilon}$ is drawn from $\mathcal{N}(0,0.1)^{S_{c}}$ and the optimal local model $\theta^{*}\in\mathbb{R}^{d}$ is drawn from the standard normal distribution where $d=11$. The training is run by FedAvg with 1 local epoch, batch size $64$, $256$, and $1024$, and 5 seeds each, respectively. The number of communication rounds are 300 and the honest-but-curious adversary only eavesdropped $d+1=12$ messages from rounds $\mathcal{T}=\{i*20|i\in\{0,1,...,11\}\}$.

### v2

@@ 6 Discussion and Conclusions
Finally, to mitigate privacy leakage, beyond differential privacy mechanisms, there are empirical defenses such as Mixup (Zhang et al. 2018), and TAPPFL (Arroyo Arevalo et al. 2024). However, the effectiveness of these defenses has not yet been demonstrated on regression tasks, and we leave this for future work.
Acknowledgements
This research was supported in part by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge, as well as by the France 2030 program, managed by the French National Research Agency under grant agreements No. ANR-21-PRRD-0005-01, ANR-23-PECL-0003, and ANR-22-PEFT-0002. It was also funded in part by the European Network of Excellence dAIEDGE under Grant Agreement Nr. 101120726, by SmartNet and LearnNet, and by the French government National Research Agency (ANR) through the UCA JEDI (ANR-15-IDEX-0001), EUR DS4H (ANR-17-EURE-0004), and the 3IA Côte d’Azur Investments in the Future project with the reference number ANR-19-P3IA-0002.
@@ A.1 Full statement and proof of Theorem 1
Consider training a least squares regression through FedAvg (Alg. 4) with batch size $B$ and local epochs $E$. Assume that
@@
    the client’s design matrix $\mathbf{x}_{c}\in\mathbb{R}^{S_{c}\times d}$ has rank $d$ equal to the number of features plus one; [^5] Remember that a client’s design matrix has a number of rows equal to the input features of the samples in the client’s local training dataset.
@@
    the components of the stochastic (mini-batch) gradient are distributed as sub-Gaussian random variables with variance proxy $\sigma^{2}$, i.e., $\mathbb{E}[\exp(\mathbf{e}(\theta)[i]/\sigma^{2})]\leq\exp(1),\forall\theta,\forall i\in\{1,2,...,d\},$ where $\mathbf{e}(\theta)=\mathbf{g}(\theta)-\nabla{\mathcal{L}_{c}}(\theta)$; [^6] Note that, this condition implies a bounded variance for each component of the stochastic gradients.
@@
    there exists $\underline{\lambda}>0$ such that $\forall n_{c}\in\mathbb{N}$, we can always select $n_{c}$ observation rounds so that $\lambda_{\min}\!\left(\frac{\Theta_{\text{out}}^{T}\Theta_{\text{out}}}{n_{c}}\right)\geq\underline{\lambda}$, where $\lambda_{\min}(\mathbf{A})$ denotes the smallest eigenvalue of the matrix $\mathbf{A}$, and $\Theta_{\text{out}}$ is defined in Alg. 2.
The error of the reconstructed model $\hat{\theta}^{*}_{c}$ of Algorithm 2 is upper bounded w.p. $\geq 1-\delta$ when $\eta\leq\frac{S_{c}}{2\lambda_{\max}(\mathbf{x}_{c}^{T}\mathbf{x}_{c})}$ and
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}=\mathcal{O}\left(\eta\sigma d\sqrt{dE\left\lceil\frac{S_{c}}{B}\right\rceil\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\cdot\underline{\lambda}}}\right).$ (7)
@@
Let $\mathbf{H}=\mathbf{x}_{c}^{T}\mathbf{x}_{c}$ which is positive definite. Let $\mathbf{y}_{c}\in\mathbb{R}^{S_{c}}$ be the labels in the local dataset ${\mathcal{D}}_{c}$ with size $S_{c}=|{\mathcal{D}}_{c}|$.
${\mathcal{L}_{c}}(\theta)=\frac{\left\lVert\mathbf{x}_{c}\theta-\mathbf{y}_{c}\right\rVert^{2}}{S_{c}}$ (8)
@@
$\mathbf{g}(\theta)=\nabla{\mathcal{L}_{c}}(\theta)+\mathbf{e}(\theta)=\frac{2}{S_{c}}\left(\mathbf{H}\theta-\mathbf{H}\theta_{c}^{*}\right)+\mathbf{e}(\theta).$ (9)
@@
$\theta_{c}^{t}(K)$ $=\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{K}\theta_{c}^{t}(0)+\left[\mathbf{I}-\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{K}\right]\theta_{c}^{*}$
$-\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{k-1}\eta\mathbf{e}_{K-k+1}.$
where $K=E\lceil\frac{S_{c}}{B}\rceil$ as in each epoch there are $\lceil\frac{S_{c}}{B}\rceil$ local steps. Let $\mathbf{W}=\left[\mathbf{I}-\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{K}\right]$. $\mathbf{W}$ is proven to be invertible in Lemma 2, then we have:
@@
$-\mathbf{W}^{-1}\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{k-1}\eta\mathbf{e}_{K-k+1}$ (10)
Let $\mathbf{\gamma}^{t}=-\eta\mathbf{W}^{-1}\sum_{k=1}^{K}\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{k-1}\mathbf{e}_{K-k+1}$. Let $\mathbf{x}[i]$ be the $i^{th}$ element of vector $\mathbf{x}$ and $\mathbf{X}[i,:]$ and $\mathbf{X}[:,i]$ be the $i^{th}$ row and column of the matrix $\mathbf{X}$, respectively. Since every element in $\mathbf{e}$ is sub-Gaussian and the weighted sum of independent sub-Gaussian variables is still sub-Gaussian, $\mathbf{\gamma}[i]$ is sub-Gaussian with $\mathbb{E}[\mathbf{\gamma}^{t}[i]]=0$ and
$\mathbb{V}\mathrm{ar}[\mathbf{\gamma}^{t}[i]]$ $\leq d\eta^{2}\sigma^{2}\lVert\mathbf{W}^{-1}\rVert_{2}^{2}\sum_{k=1}^{K}\left\lVert\left(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{k-1}\right\rVert^{2}_{2}$
$=\frac{d\eta^{2}\sigma^{2}}{\lambda^{2}_{\text{min}}(\mathbf{W})}\sum_{k=1}^{K}\left\lVert(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H})^{k-1}\right\rVert^{2}_{2}$
$=\frac{d\eta^{2}\sigma^{2}}{\lambda^{2}_{\text{min}}(\mathbf{W})}\sum_{k=1}^{K}\rho_{\max}^{2}((\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H})^{k-1}),$
@@
Let $\Lambda$ be a diagonal matrix whose entries are the positive eigenvalues of $\mathbf{H}$ and $\lambda_{\max}(\mathbf{H})$ be the largest eigenvalue of $\mathbf{H}$, i.e., $\lambda_{\max}(\mathbf{H})=\lVert\Lambda\rVert_{1}$. When $\eta\leq\frac{S_{c}}{2\lambda_{\max}(\mathbf{H})}$, the diagonal values in $\frac{2\eta}{S_{c}}\Lambda$ are positive and smaller than or equal to $1$. Moreover, the matrix $(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H})^{k-1}$ is positive semi-definite and we have $\rho_{\max}[(\mathbf{I}-\frac{2\eta}{S_{c}}\mathbf{H})^{k-1}]\leq 1$. Thus, we have
@@
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}^{2}\leq\frac{d^{3}\eta^{2}\sigma^{2}K}{\lambda_{\text{min}}^{2}(\mathbf{W})}\cdot\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\underline{\lambda}},\;\;\text{w.p.}\geq 1-\delta.$ (18)
@@
$\mathbf{W}$ $=I-\left(I-\frac{2\eta}{S_{c}}\mathbf{H}\right)^{E\left\lceil\frac{S_{c}}{b}\right\rceil}$
$=\mathbf{U}\left(I-\left(I-\frac{2\eta}{S_{c}}\Lambda\right)^{E\left\lceil\frac{S_{c}}{b}\right\rceil}\right)\mathbf{U}^{T}.$
@@
$\lambda_{\min}(\mathbf{W})$ $=1-\left(1-\frac{2\eta}{S_{c}}\lambda_{\min}(\mathbf{H})\right)^{E\left\lceil\frac{S_{c}}{b}\right\rceil}$
$\lambda_{\max}(\mathbf{W})$ $=1-\left(1-\frac{2\eta}{S_{c}}\lambda_{\max}(\mathbf{H})\right)^{E\left\lceil\frac{S_{c}}{b}\right\rceil}$
@@ A.2 Proof of Proposition 2
$\mathbf{g}(\theta)=\nabla{\mathcal{L}_{c}}(\theta)=\frac{2}{S_{c}}\left(\mathbf{H}\theta-\mathbf{H}\theta_{c}^{*}\right)$
@@ A.3 Proof of Proposition 3
${\mathcal{L}_{\mathbf{c}}}(\theta)$ $=$ $\frac{1}{S_{c}}\left(\theta^{T}\mathbf{H}_{\mathbf{c}}\theta-2(\mathbf{x}_{c}^{T}\mathbf{y}_{c})^{T}\theta-\mathbf{y}_{c}^{T}\mathbf{y}_{c}\right)$ (20)
$=$ $\frac{1}{S_{c}}\left(\sum_{i=1}^{d}\theta_{[i]}^{2}-\sum_{i=1}^{d-1}\theta_{[i]}\theta_{[i+1]}-\theta_{[1]}-C\right)$
@@
Let $n=|{\mathcal{C}}|$ be the number of clients and assume that every client has $S_{c}$ local data samples. Then the global empirical risk and its gradients have the following expressions:
${\mathcal{L}}(\theta)$ $=\frac{1}{nS_{c}}\left(\sum_{i=1}^{d}n\times\theta_{[i]}^{2}-\sum_{i=1}^{d-1}\theta_{[i]}\theta_{[i+1]}-\theta_{[1]}-C^{\prime}\right),$ (21)
$\nabla{\mathcal{L}}(\theta)_{[1]}$ $=\frac{1}{nS_{c}}(2n\times\theta_{[1]}-\theta_{[2]}-1),$ (22)
$\nabla{\mathcal{L}}(\theta)_{[i]}$ $=\frac{1}{nS_{c}}\left(2n\times\theta_{[i]}-\theta_{[i-1]}-\theta_{[i+1]}\right),\forall i\in\{2,...,d-1\},$ (23)
$\nabla{\mathcal{L}}(\theta)_{[d]}$ $=\frac{1}{nS_{c}}\left(2n\times\theta_{[d]}-\theta_{[d-1]}\right),$ (24)
@@ Appendix B The details of experimental setups
B.1 Client’s update rule
In all our experiments we assume each client follows FedAvg update rule, performing locally multiple stochastic gradient steps. The detailed update rule is in Alg 4.
2:  for each local epoch $e$ from 1 to $E$ do
3:     for batch $b\in\mathcal{B}$ do
5:        $k\leftarrow k+1$
6:  Return $\theta_{c}^{t}(K)$, where $K=E\lceil\frac{S_{c}}{B}\rceil$
B.2 Experiment in Figure 1
We evaluate different AIAs when four clients train a neural network (a single hidden layer of 128 neurons using ReLU as activation function) through FedAvg with 1 local epoch and batch size 32. Each client stores $S_{c}=|{\mathcal{D}}_{c}|$ data points randomly sampled from ACS Income dataset (Ding et al. 2024). This dataset contains census information from 50 U.S. states and Puerto Rico, spanning from 2014 to 2018. It includes 15 features related to demographic information such as age, occupation, and education level. The adversary infers the gender attribute of every data sample held by the client given access to the released (public) information.
@@ B.2 Experiment in Figure 1
B.3 Toy dataset setting
We test our proposed passive LMRA (Alg. 2) on a toy federated least squares regression with two clients. Each client $c$ has 1024 samples with 10 features, where nine numerical features are sampled from a uniform distribution over $[0,1)$ and one binary feature is sampled from a Bernoulli distribution with $p=1/2$. The prediction $\mathbf{y}$ is generated from the regression model $\mathbf{y}=\mathbf{x}_{c}\theta^{*}+\mathbf{\epsilon}$ where $\mathbf{\epsilon}$ is drawn from $\mathcal{N}(0,0.1)^{S_{c}}$ and the optimal local model $\theta^{*}\in\mathbb{R}^{d}$ is drawn from the standard normal distribution where $d=11$. The training is run by FedAvg with 1 local epoch, batch size $64$, $256$, and $1024$, and 5 seeds each, respectively. The number of communication rounds are 300 and the honest-but-curious adversary only eavesdropped $d+1=12$ messages from rounds $\mathcal{T}=\{i*20|i\in\{0,1,...,11\}\}$.